# VP-MEL: Visual Prompts Guided Multimodal Entity Linking

### v3

@@ Abstract
Multimodal entity linking (MEL), a task aimed at linking mentions within multimodal contexts to their corresponding entities in a knowledge base (KB), has attracted much attention due to its wide applications in recent years. However, existing MEL methods often rely heavily on mention words as retrieval cues, which limits their ability to effectively utilize information from both images and text. This reliance poses significant challenges in scenarios where mention words are absent, as current MEL approaches struggle to leverage image-text pairs for accurate entity linking. To solve these issues, we introduce a Visual Prompts guided Multimodal Entity Linking (VP-MEL) task. Given a text-image pair, VP-MEL aims to link a marked region (i.e., visual prompt) in an image to its corresponding entities in the knowledge base. To facilitate this task, we present a new dataset, VPWiki, specifically designed for VP-MEL. Furthermore, we propose a framework named FBMEL, which enhances visual feature extraction using visual prompts and leverages the pretrained Detective-VLM model to capture latent information. Experimental results on the VPWiki dataset demonstrate that FBMEL outperforms baseline methods across multiple benchmarks for the VP-MEL task. Code and datasets will be released at https://anonymous.4open.science/r/VP-MEL6-F9B0.
@@ 1 Introduction
Figure 1: Comparison between MEL and VP-MEL tasks. MEL struggles with entity linking for image content when mention words are unavailable. In contrast, VP-MEL addresses this limitation by using visual prompts to link specific regions in the images to the correct entities in the knowledge base.
@@
Text information dependency: MEL heavily relies on mention words for entity linking, which often exhibit substantial overlap with entity names in practical applications. This inherent similarity enables mention words to serve as strong cues for identifying specific entities within the knowledge base (KB). Although some ambiguity remains, this overlap substantially simplifies the entity linking process. However, MEL struggles to perform effectively and often produces incorrect outputs when mention words are unavailable. As shown in Figure 1, the expected mention is Ted Kennedy (on the right), but MEL generates incorrect responses, like Lolo Soetoro and Hartford. Many results, such as Lolo Soetoro, are entirely unrelated to both the image and text. This excessive reliance on mention words confines MEL to a text-focused task. It is necessary to introduce a new MEL task to overcome such constraints, allowing MEL to broaden its application scenarios and adapt to the multimodal era.
Image modality impurity: Compared to the text modality, the image modality often contains more noise. As illustrated in Figure 1, only the person on the right is relevant, while the one on the left represents noise. Misinterpreting or misusing such image information can seriously affect the results. Most existing coarse-grained methods (Yang et al., 2023; Luo et al., 2023; Wang et al., 2022a) encode the entire image directly, making it difficult to eliminate noise interference. Song et al. (2024) attempts to enhance visual features by utilizing fine-grained image attributes, but they heavily depends on high-quality images and often increase model size and training costs. To better leverage image information, it is necessary for MEL to find a way to overcome noise interference in images and focus on the regions related to the mentions.
Existing methods demonstrate limited effectiveness in utilizing image information and are overly dependent on mention words as textual cues. In many application scenarios, the challenge of MEL is considerably worsened by the inability of users to provide relevant mention words. So we ask: Is it possible to introduce a MEL task that does not depend on mention words, thereby avoiding excessive reliance on text and enhancing the performance of image-based reasoning?
@@
In this paper, we define a new Visual Prompts guided Multimodal Entity Linking (VP-MEL) task. As shown in Figure 1, VP-MEL is designed to address the entity linking task for image-text pairs. VP-MEL employs a straightforward method to annotate mentions directly onto images using visual prompts, eliminating the need for mention words or any additional complex operations. Free from the constraints of mention words, VP-MEL supports a broader range of applications. It enables effective entity linking even in situations where users are unfamiliar with the textual language or place greater emphasis on image-based information. To support this new task, we construct a VPWiki dataset based on the existing MEL public datasets, where visual prompts are annotated for each mention within the associated images.
We then propose a Feature Balanced Multimodal Entity Linking (FBMEL) framework. Extensive research (Cai et al., 2024; Shtedritski et al., 2023) demonstrates that pre-trained CLIP visual encoders can effectively interpret visual markers. In FBMEL, the CLIP visual encoder is employed to extract both global image features and local features guided by visual prompts. Additionally, a Vision-Language Model (VLM) equipped with CLIP visual encoder is pre-trained to generate useful text information from visual prompts, serving as an additional supplement to the text information. FBMEL enables visual prompts to deliver both supplementary visual and textual information, ensuring a balanced utilization of image and text while avoiding excessive reliance on a single modality.
@@
    We introduce VP-MEL, a new entity linking task that replaces traditional mention words with visual prompts to annotate mentions directly in the images.
@@
    We propose the FBMEL framework to address VP-MEL task by effectively leveraging multimodal information and reducing reliance on a single modality. Compared to prior methods, FBMEL achieves a 20% performance improvement in the VP-MEL task and maintains competitive results in the MEL task.
@@ 2.1 Multimodal Entity Linking
Given the widespread use of image-text content in social media, the integration of both modalities for entity linking is essential. For example, Moon et al. (2018) pioneer the use of images to aid entity linking. Building on this, Adjali et al. (2020) and Gan et al. (2021) construct MEL datasets from Twitter and long movie reviews. Expanding the scope of MEL datasets, Wang et al. (2022c) present a high-quality MEL dataset from Wikinews, featuring diversified contextual topics and entity types. To achieve better performance on these datasets, a multitude of outstanding works in the MEL field (Wang et al., 2022a; Yang et al., 2023; Luo et al., 2023; Shi et al., 2024; Song et al., 2024) emerge, focusing on extracting and interacting with multimodal information. Although multimodal information can enhance entity linking performance, in these methods, text consistently dominates over images.
@@ 3 Dataset
To improve data annotation efficiency, we develop a pipeline that automatically annotates visual prompts in images based on mention words inspired by Li et al. (2024). In the pipeline, the Visual Entailment Module is employed to evaluate and filter out the highly relevant data. Subsequently, the Visual Grounding Module annotates the visual prompts in the images. The details of modules in the pipeline are provided in the Appendix A.6. The annotation team consists of 10 annotators and 2 experienced experts. All annotators have linguistic knowledge and are instructed with detailed annotation principles. Fleiss Kappa score (Fleiss, 1971) of annotators is 0.83, indicating strong agreement among them. We use the Intersection over Union (IoU) metric to assess annotation quality and discard samples with an IoU score below 0.5.
Figure 2: An example from VPWiki and WikiDiverse. The red box in the left image represents the visual prompt annotated for the VP-MEL task. The red text in the right image shows the annotated mention words.
@@
Figure 2 illustrates an example from the VPWiki dataset and highlights its differences from traditional MEL data. The VPWiki dataset comprises a total of 12,720 samples, which are randomly split into training, validation, and test sets with an 8:1:1 ratio. Detailed statistics of the VPWiki dataset are provided in Table 1. Additionally, Figure 3 presents the distribution of entity types and the number of candidate entities per mention in the dataset. In Figure 3(a), abbreviations are used to represent each entity type. Meanwhile, Figure 3(b) shows that as the number of candidate entities per mention increases, the task becomes increasingly challenging.
@@ 5 Methodology
In this section, we describe the proposed FBMEL framework for the VP-MEL task. As illustrated in Figure 4, FBMEL utilizes visual encoder to extract both deep semantic features and shallow texture features, which are enhanced by visual prompts (§5.1). To avoid excessive reliance on visual features, the Detective-VLM module is designed to generate supplementary textual information guided by visual prompts(§5.2), which is then combined with the original text and processed by the text encoder (§5.3). Finally, a similarity score is computed after integrating the visual and textual features(§5.4).
Figure 4: The overall architecture of Feature Balanced Multimodal Entity Linking (FBMEL) framework. The image-text pairs of the Mention and Entity are used together as input. Specifically, Mention Text is the sentence corresponding to Mention Image, while Entity Text consists of Entity Name and Entity Attribute corresponding to the Entity Image in the Knowledge Base.
@@ 5.1 Visual Encoder
To extract image features, we choose pre-trained CLIP model(Dosovitskiy et al., 2021) as our visual encoder. The image $m_{v_{j}}$ of $M_{j}$ is reshaped into n 2D patches. After this, image patches are processed through visual encoder to extract features. The hidden states extracted from $m_{v_{j}}$ by the CLIP visual encoder are represented as $V_{M_{j}}^{l}=\left[v_{[CLS]}^{0};v_{M_{j}}^{1};v_{M_{j}}^{2};...;v_{M_{j}}^{n}\right]\in\mathbb{R}^{(n+1)\times d_{c}}$, where $d_{c}$ denotes the dimension of the hidden state and $l$ denotes the number of layers in the encoder.
@@ 5.2 Detective-VLM
To prevent the model from becoming overly dependent on visual information, it is necessary to extract additional textual information to balance contributions of both modalities.
Most VLMs (Liu et al., 2024; Zhu et al., 2024; Ye et al., 2023; Li et al., 2023) choose to use the CLIP visual encoder, which means they have a better ability to focus on the markers in images compared to other visual methods (Cai et al., 2024; Shtedritski et al., 2023). Therefore, we instruction fine-tune a VLM, designed to extract effective information from images. The VLM follows template designed below to further mine potential information from image $m_{v_{j}}$ and sentence $m_{s_{j}}$ of mention $M_{j}$, assisting in subsequent feature extraction:
Background: $\{Image\}$ Text: $\{Sentence\}$ Question: In the red box of the image, tell me briefly what is the $\{Entity\ \allowbreak Type\}$, and which $\{Entity\ \allowbreak Name\}$ in the text corresponds to the $\{Entity\ \allowbreak Type\}$? Answer: $\{Entity\ \allowbreak Name\}$$\{Entity\ \allowbreak Type\}$
@@ 5.5 Contrastive Learning
[…]
@@ 6.1 Experimental Settings
All the training and testing are conducted on a device equipped with 4 Intel(R) Xeon(R) Platinum 8380 CPUs and 8 NVIDIA A800-SXM4-80GB GPUs. Detailed experimental settings are provided in Appendix A.1. To comprehensively evaluate the effectiveness of our approach, we compare FBMEL with various competitive MEL baselines and VLM baselines. A detailed introduction of these baselines is provided in the Appendix A.2.
@@ 6.2 Main Results
As shown in Table 2(a), FBMEL significantly outperforms all other methods on VP-MEL task. First, among the VLM methods, LLaVA-1.5 has the smallest performance gap compared to our method, with differences of 5.14%, 1.97%, and 10.58% from FBMEL across the three metrics, respectively. Even so, considering the substantial data and computational resources required to train LLaVA, FBMEL still demonstrates an obvious advantage. Second, there is a notable performance gap between MEL methods and FBMEL. MEL methods struggle with effective entity linking in scenarios where mention words are absent, underscoring their limitations and the robustness of our approach.
@@
Table 2(b) presents the experimental results comparing FBMEL with other methods on MEL dataset. During the testing process, the Detective-VLM is removed from FBMEL, and its output is replaced with mention words from WikiDiverse. The performance of FBMEL is slightly behind MIMIC and MELOV, with differences of only 0.27%, 0.28%, and 0.07% across the three metrics. This is expected, as FBMEL places greater emphasis on image features while reducing reliance on text features. Furthermore, the absence of visual prompts in the images limits the ability of FBMEL to fully utilize visual information.
@@
FBMEL 63.24 79.76 86.36 23.87 38.37 45.14
[…]
@@ 6.3 Detailed Analysis
For a fair comparison, we remove the Detective-VLM module from FBMEL. As shown in Table 3, the performance of MEL methods drop significantly across all three metrics in the absence of mention words. The average performance decline is 72.65%, 64.48%, and 60.28%, respectively. This indicates that MEL methods fail to extract meaningful information from visual and textual data, making them unsuitable for tasks without mention words. In contrast, even without Detective-VLM, visual prompts, or mention words, FBMEL can still achieve the best metrics. This demonstrates that FBMEL in the VP-MEL task possesses a stronger capability to leverage both image and text information effectively.
@@
As shown in Table 4, we evaluate the effectiveness of Detective-VLM by replacing it with various VLMs and analyzing the results. Our method achieves the best performance across all metrics. In particular, Detective-VLM shows an absolute improvement of 5.89% in Hit@1 compared to the second-best approach. In contrast, non-fine-tuned VLMs often produce a large amount of irrelevant information, which hampers subsequent processing.
@@
FBMEL 48.34 67.53 77.50 82.63 87.92
FBMEL 35.65 53.93 65.26 73.57 80.51
FBMEL 35.03 53.80 65.01 73.26 80.39
@@
In Table 6, we conduct ablation study on the FBMEL framework. First, we remove the Detective-VLM module from FBMEL, which results in a decline across all metrics. Notably, even without Detective-VLM, FBMEL shows robust entity linking performance, outperforming MEL methods as shown in Table 2(a). This highlights the ability of FBMEL to efficiently leverage multimodal information from both images and text. To further evaluate the impact of Visual Prompts, we remove the visual prompts from the images in the previous setup. This leads to a noticeable drop in all metrics, underscoring the importance of visual prompts in guiding the model to focus on relevant regions within images.

### v4

@@ Abstract
Multimodal entity linking (MEL), a task aimed at linking mentions within multimodal contexts to their corresponding entities in a knowledge base (KB), has attracted much attention due to its wide applications. However, existing MEL methods primarily rely on mention words as retrieval cues, which limits their ability to effectively utilize both textual and visual information. As a result, MEL struggles to retrieve entities accurately, particularly when the focus is on image objects or when mention words are absent from the text. To solve these issues, we introduce Visual Prompts guided Multimodal Entity Linking (VP-MEL). Given a text-image pair, VP-MEL links a marked image region (i.e., visual prompt) to its corresponding KB entity. To support this task, we construct VPWiki, a dataset specifically designed for VP-MEL. Additionally, we propose the Implicit Information-Enhanced Reasoning (IIER) framework, which enhances visual feature extraction through visual prompts and leverages the pretrained Detective-VLM model to capture latent information. Experimental results on VPWiki demonstrate that IIER outperforms baseline methods across multiple benchmarks for VP-MEL.
@@ 1 Introduction
Figure 1: Comparison between MEL and VP-MEL tasks. MEL is typically limited to selecting mentions from text. In contrast, VP-MEL addresses this limitation by using visual prompts to link specific regions in the images to the correct entities in the knowledge base.
@@
Text information dependency: MEL relies on mention words for entity linking, as these words often overlap significantly with entity names in real-world applications. This overlap provides strong cues for identifying entities within the knowledge base (KB). However, MEL performs poorly when mention words are absent or unannotated. As shown in Figure 1, without annotated mention words, MEL computes similarity based on the entire text, which can lead to erroneous linking. For instance, it might associate the data with the entity Olympic Games due to high textual similarity. Since Lee C W and Lin Dan are not explicitly mentioned, MEL fails to establish correct links to these entities. This issue underscores the difficulty of MEL in accurately linking data when mention words related to the entities are missing from the data.
Image modality impurity: Compared to text, images often contain more noise. Misinterpreting or misusing image information can significantly affect the results. Most existing coarse-grained methods (Yang et al., 2023; Luo et al., 2023; Wang et al., 2022a) directly encode the entire image, making it difficult to eliminate noise interference. Song et al. (2024) improves MEL performance by extracting fine-grained image information through object detection. However, this approach still relies on sufficient textual information for accurate object localization and is prone to interference from visually similar objects. Therefore, a potentially effective strategy to mitigate image noise is enhancing object localization precision while reducing dependence on textual data.
These limitations hinder the ability of MEL to leverage multimodal data effectively. Images, despite being fundamental to multimodal data, contribute minimally to MEL. Moreover, reliance on textual data constrains MEL performance when text is scarce or incomplete. So we ask: Is it possible to link specific objects in multimodal images to the KB even with insufficient textual information? Investigating this possibility could unlock the full potential of multimodal data in MEL.
@@
In this paper, we introduce Visual Prompts guided Multimodal Entity Linking (VP-MEL), a new task designed for entity linking in image-text pairs, as shown in Figure 1. VP-MEL annotates mentions directly on images using visual prompts, eliminating reliance on textual mention words. This approach expands the application scope, enabling effective linking of multimodal data to the KB even when textual information is limited or image-based information is prioritized. To support this task, we construct the VPWiki dataset based on existing MEL public datasets, where visual prompts are annotated for each mention within the associated images.
To address VP-MEL, we propose the Implicit Information-Enhanced Reasoning (IIER) framework. IIER leverages visual prompts as guiding texture features to focus on specific local image regions. To reduce reliance on textual data, it utilizes an external implicit knowledge base to heuristically generate auxiliary information for the reasoning process. Specifically, a CLIP visual encoder is employed to extract both global image features and local features guided by visual prompts. AdditionlLy, a Vision-Language Model (VLM) equipped with CLIP visual encoder is pre-trained to generate textual information from visual prompts, supplementing existing text data. IIER provides both supplementary visual and textual information, effectively linking objects in images to the KB.
@@
    We introduce VP-MEL, a new entity linking task that replaces traditional mention words with visual prompts, linking specific objects in images to the KB.
@@
    We propose the IIER framework to address VP-MEL task by effectively leveraging multimodal information and reducing reliance on a single modality. Compared to prior methods, IIER achieves a 20% performance improvement in the VP-MEL task and maintains competitive results in the MEL task.
@@ 2.1 Multimodal Entity Linking
Given the widespread use of image-text content in social media, the integration of both modalities for entity linking is essential. For example, Moon et al. (2018) pioneer the use of images to aid entity linking. Building on this, Adjali et al. (2020) and Gan et al. (2021) construct MEL datasets from Twitter and long movie reviews. Expanding the scope of MEL datasets, Wang et al. (2022c) present a high-quality MEL dataset from Wikinews, featuring diversified contextual topics and entity types. To achieve better performance on these datasets, a multitude of outstanding works in the MEL field (Wang et al., 2022a; Yang et al., 2023; Luo et al., 2023; Shi et al., 2024) emerge, focusing on extracting and interacting with multimodal information. Song et al. (2024) use object detection to extract visual information from images and better link mention words to correct entities, but still face difficulty in linking images to KBs in the absence of mention words. Although multimodal information can enhance entity linking performance, in these methods, text consistently dominates over images.
@@ 3 Dataset
To improve data annotation efficiency, we develop a pipeline that automatically annotates visual prompts in images based on mention words inspired by Li et al. (2024). In the pipeline, the Visual Entailment Module is employed to evaluate and filter out the highly relevant data. Subsequently, the Visual Grounding Module annotates the visual prompts in the images. The details of the pipeline are provided in Appendix A.6. The annotation team consists of 10 annotators and 2 experienced experts. All annotators have linguistic knowledge and are instructed with detailed annotation principles. Fleiss Kappa score (Fleiss, 1971) of annotators is 0.83, indicating strong agreement among them. We use the Intersection over Union (IoU) metric to assess annotation quality and discard samples with an IoU score below 0.5.
Figure 2: An example from VPWiki. GT denotes the ground truth entity. The red box in the left image represents the visual prompt annotated for the VP-MEL task.
@@
Figure 2 illustrates an example from the VPWiki dataset and highlights its differences from traditional MEL data. Additional data samples are provided in A.8. The VPWiki dataset comprises a total of 12,720 samples, which are randomly split into training, validation, and test sets with an 8:1:1 ratio. Detailed statistics of the VPWiki dataset are provided in Table 1. Additionally, Figure 3 presents the distribution of entity types and the number of candidate entities per mention in the dataset. In Figure 3(a), abbreviations are used to represent each entity type. Meanwhile, Figure 3(b) shows that as the number of candidate entities per mention increases, the task becomes increasingly challenging.
@@ 5 Methodology
In this section, we describe the proposed IIER framework for the VP-MEL task. As illustrated in Figure 4, IIER utilizes visual encoder to extract both deep semantic features and shallow texture features, which are enhanced by visual prompts (§5.1). To avoid excessive reliance on visual features, the Detective-VLM module is designed to generate supplementary textual information guided by visual prompts(§5.2), which is then combined with the original text and processed by the text encoder (§5.3). Finally, a similarity score is computed after integrating the visual and textual features(§5.4).
Figure 4: The overall architecture of Implicit Information-Enhanced Reasoning (IIER) framework. The image-text pairs of the Mention and Entity are used together as input. Specifically, Mention Text is the sentence corresponding to Mention Image, while Entity Text consists of Entity Name and Entity Attribute corresponding to the Entity Image in the Knowledge Base.
@@ 5.1 Visual Encoder
We choose pre-trained CLIP model (Dosovitskiy et al., 2021) as our visual encoder. Extensive research (Cai et al., 2024; Shtedritski et al., 2023) demonstrates its effectiveness in interpreting visual markers. The image $m_{v_{j}}$ of $M_{j}$ is reshaped into n 2D patches. After this, image patches are processed through visual encoder to extract features. The hidden states extracted from $m_{v_{j}}$ by the CLIP visual encoder are represented as $V_{M_{j}}^{l}=\left[v_{[CLS]}^{0};v_{M_{j}}^{1};v_{M_{j}}^{2};...;v_{M_{j}}^{n}\right]\in\mathbb{R}^{(n+1)\times d_{c}}$, where $d_{c}$ denotes the dimension of the hidden state and $l$ denotes the number of layers in the encoder.
@@ 5.2 Detective-VLM
Real-world multimodal data often contain challenges such as short texts or image noise. In this context, VLMs serve as implicit knowledge bases, can analyze both image and text to infer useful auxiliary information. Most VLMs (Liu et al., 2024; Zhu et al., 2024; Ye et al., 2023; Li et al., 2023) adopt the CLIP visual encoder, enabling them to focus more effectively on markers in images compared to other visual methods (Cai et al., 2024; Shtedritski et al., 2023). Therefore, we instruction fine-tune a VLM to extract effective information from images. The VLM follows template designed below to further mine potential information from the image $m_{v_{j}}$ and sentence $m_{s_{j}}$ of mention $M_{j}$, assisting in subsequent feature extraction:
Background: $\{Image\}$ Text: $\{Sentence\}$ Question: Based on the text ’$\{Sentence\}$’, tell me briefly what is the $\{Entity\ \allowbreak Type\}$ and $\{Entity\ \allowbreak Name\}$ in the red box of the $\{Image\}$? Answer: $\{Entity\ \allowbreak Name\}$$\{Entity\ \allowbreak Type\}$
@@ 5.5 Contrastive Learning
Table 2: Performance comparison on the VP-MEL(a) and MEL(b) tasks. Baseline results marked with "*" are based on Sui et al. (2024). Each method is run 5 times with different random seeds, and the mean value of each metric is reported. The best score is highlighted in bold. Detailed evaluation metrics can be found in Appendix A.3.
@@ 6.1 Experimental Settings
All the training and testing are conducted on a device equipped with 4 Intel(R) Xeon(R) Platinum 8380 CPUs and 8 NVIDIA A800-SXM4-80GB GPUs. Detailed experimental settings are provided in Appendix A.1. To comprehensively evaluate the effectiveness of our approach, we compare IIER with various competitive MEL baselines and VLM baselines. A detailed introduction of these baselines is provided in the Appendix A.2.
@@ 6.2 Main Results
As shown in Table 2(a), IIER significantly outperforms all other methods on VP-MEL task. First, among the VLM methods, LLaVA-1.5 has the smallest performance gap compared to our method, with differences of 5.16%, 3.16%, and 11.79% from IIER across the three metrics, respectively. Even so, given the significantly lower training cost compared to LLaVA, IIER offers a clear advantage in efficiency while achieving competitive performance. Second, there is a notable performance gap between MEL methods and IIER. MEL methods struggle with effective entity linking in scenarios where mention words are absent, underscoring their limitations and the robustness of our approach.
@@
Table 2(b) presents the experimental results comparing IIER with other methods on MEL dataset. During testing, the Detective-VLM analyzes image and text data to generate a concise representation of mention words, which are concatenated with the text and used for entity linking similarity calculation. With enhanced visual features and external knowledge, IIER demonstrates excellent performance in the MEL task. Although our work primarily focuses on VP-MEL rather than MEL, IIER still demonstrates strong competitiveness compared to the state-of-the-art MEL method. This highlights the effectiveness of external implicit knowledge in supporting the reasoning process of entity linking.
@@
IIER 69.47 84.43 88.79 23.87 38.37 45.14
Table 3: Performance comparison in the absence of mention words on the WikiDiverse dataset. The symbol "*" represents the dataset without annotated mention words.
@@ 6.3 Detailed Analysis
As shown in Table 3, the performance of MEL methods drop significantly across all three metrics in the absence of mention words. The average performance decline is 72.65%, 64.48%, and 60.28%, respectively. This indicates that MEL methods fail to extract meaningful information from visual and textual data, making them unsuitable for tasks without mention words. In contrast, even without Detective-VLM, visual prompts, or mention words, IIER can still achieve the best metrics. This demonstrates that IIER in the VP-MEL task possesses a stronger capability to leverage both image and text information effectively.
@@
As shown in Table 4, we evaluate the effectiveness of Detective-VLM by replacing it with various VLMs and analyzing the results. Our method achieves the best performance across all metrics. In particular, Detective-VLM shows an absolute improvement of 5.91% in Hit@1 compared to the second-best approach. In contrast, non-fine-tuned VLMs often produce a large amount of irrelevant information, which hampers subsequent processing.
@@
IIER 48.36 67.51 77.50 82.59 87.90
IIER 35.65 53.93 65.26 73.57 80.51
IIER 35.03 53.80 65.01 73.26 80.39
@@
In Table 6, we conduct ablation study on the IIER framework. First, we remove the Detective-VLM module from IIER, which results in a decline across all metrics. Notably, even without Detective-VLM, IIER shows robust entity linking performance, outperforming MEL methods as shown in Table 2(a). This highlights the ability of IIER to efficiently leverage multimodal information from both images and text. Subsequently, removing the visual prompts from the images results in a decline across all metrics, emphasizing the crucial role of visual prompts in guiding the model to focus on relevant regions within the images. Note that the slight decrease in metrics does not suggest a diminished significance of visual prompts, as they are integral to the functioning of the VLM.